OpenStudy (anonymous):

When solving negative one over eight (x + 35) = −7, what is the correct sequence of operations? Multiply each side by negative one over eight , add 35 to each side Multiply each side by negative one over eight , subtract 35 from each side Multiply each side by −8, subtract 35 from each side Multiply each side by −8, add 35 to each side

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: If I start with this equation: \[\Large - \frac{{x + 35}}{8} = - \frac{1}{7}\] then I can multiply both sides by -8, so I can write this: \[\Large x + 35 = \frac{8}{7}\]

OpenStudy (anonymous):

c

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

;)

OpenStudy (michele_laino):

:)